# Are Neuromorphic Architectures Inherently Privacy-preserving? An Exploratory Study

### v1

@@ 4. Methodology
This study aims to assess whether SNNs demonstrate an advantage in privacy resilience compared to ANNs. The methodology is divided into three experimental phases: (1) evaluating privacy risk under MIAs for both ANNs and SNNs, (2) investigating SNN privacy vulnerability across various algorithms, and (3) examining the privacy-utility tradeoff using Differentially Private Stochastic Gradient Descent (DPSGD).
These phases are outlined in the diagram (Figure 1) to illustrate the workflow and interconnections between the models and methods used.
@@ 4.1. Comparison of Privacy Vulnerability between ANNs and SNNs
To compare privacy risks, we applied MIAs to both ANNs and SNNs across equivalent architectures: convolutional baseline models, ResNet18, and VGG16. In this phase, shadow models were trained on labeled datasets (Figure 1, top-right) to replicate the behavior of the target model, and attack models were trained to determine the membership of specific data points in the training set. The effectiveness of these attacks was measured using the Area Under the ROC Curve (AUC) for both ANN and SNN models. This comparison allows for a direct analysis of the vulnerability of each model type to MIAs (Figure 1, yellow block).
@@ 4.2. Algorithmic Exploration within the SNN Architecture
The second phase focused on evaluating SNNs under different algorithmic conditions to assess the consistency of their privacy resilience. Three frameworks were explored: Surrogate Gradient Algorithm, Evolutionary Algorithm, and Intel’s LAVA framework.
Surrogate Gradient Algorithm: The baseline SNN model was tested with three different encoders (Delta, Latency, Delta Modulation) from the snntorch library (Figure 1, left section, "SNN Exploration Space"). The goal was to analyze the impact of these encoding mechanisms on the model’s susceptibility to MIAs by examining variations in spike generation methods and their effect on attack success rates.
Evolutionary Algorithm: Using the TennLab framework, we systematically varied parameters such as Population Size, Crossover Rate, Mutation Rate, Processor configurations, and Encoder settings. These variations were implemented to determine their impact on the privacy vulnerability of SNN models (Figure 1, green block).
@@ 4.3. Privacy-Utility Trade-off Analysis
In the final phase of our experiments, we applied DPSGD to both ANN and SNN models (Figure 1, bottom) to evaluate the privacy-utility tradeoff. DPSGD introduces noise into the gradient updates during training, providing differential privacy protection at the expense of model accuracy. We implemented DPSGD on the convolutional, ResNet18, and VGG16 architectures to quantify the impact of privacy measures on model performance (accuracy) and privacy (attack AUC). This allowed for a comparative analysis of the privacy-utility balance between ANNs and SNNs under identical privacy constraints (Figure 1, blue block).
@@ 5.1. Dataset and Model Architecture
Network Variant Structure Params*
@@
We evaluate our method on both image and tabular datasets. For image classification tasks, we use MNIST (LeCun et al., 2010) and Fashion-MNIST (Xiao et al., 2017) (28×28 grayscale images, 10 classes), CIFAR-10 (Recht et al., 2018) and CIFAR-100 (Sharma et al., 2018) (32×32 RGB images, 10 and 100 classes). We also include two tabular datasets: Iris (Omelina et al., 2021) (4 features, 3 classes) and Breast Cancer  (Zwitter and Soklic, 1988) (30 features, 2 classes).
As shown in Table  1, we implement three model architectures in both ANN and SNN configurations. The baseline model uses two convolutional layers for image data and a fully connected architecture for tabular data. ResNet18 consists of four basic blocks with group normalization and skip connections. VGG16 uses 13 convolutional layers with filter sizes increasing from 64 to 512, followed by three fully connected layers. For the SNN variants, we replace ReLU activations with leaky integrate-and-fire (LIF) neurons and add temporal processing components. We implement all ANN models using PyTorch (Ketkar et al., 2021). For SNN implementations, we use snnTorch for the baseline model and SpikingJelly (Fang et al., 2023) for ResNet and VGG16 architectures.
@@ 5.2. MIA
We implement the MIA framework using a two-model approach as illustrated in Figure 2. The experimental setup includes a ’target’ model and a ’shadow’ model, where we evaluate three different architectures: baseline model, ResNet18, and VGG16. The shadow model simulates target’s functionality by training on 80% of the same data, maintaining identical architectural framework as the target model. For the attack phase, we employ a Support Vector Machine (SVM) as our attack model, chosen for its effectiveness as a binary classifier to determine whether a data point is a member or non-member of the training set. We evaluate the models by querying with their respective training and test sets, labeling the responses as ’IN’ if the data was in the model’s training set and ’OUT’ otherwise. To prevent the classifier’s bias towards the more frequent classes, we implement undersampling of the predominant class. The SVM attack model, trained on these labeled predictions from the shadow model, provides a quantified measure of the model’s susceptibility to MIA through its ability to accurately classify membership status.
@@ 5.3. DPSGD
We implement DPSGD as shown in Figure 3 using the "Privacy Engine" module from Opacus (Yousefpour et al., 2021). The DPSGD algorithm  1 begins by selecting a minibatch $L_{t}$ uniformly from the set of indices $\{1,\dots,N\}$ with size $L$. For each data point $x_{i}$ in the minibatch, the gradient of the loss function $\mathcal{L}(\theta_{t},x_{i})$ with respect to the parameters $\theta_{t}$ is computed as $g_{t}(x_{i})$. These gradients are then clipped using an $L2$ norm operation, $\hat{g}_{t}(x_{i})=g_{t}(x_{i})/\max(1,\|g_{t}(x_{i})\|_{2}/C)$. Gaussian noise $\mathcal{N}(0,\sigma^{2}C^{2}I)$ is added to the average of the clipped gradients, with $\sigma$ being the noise scale, to ensure differential privacy. The noisy gradient $\tilde{g}_{t}$ is used to update the model parameters: $\theta_{t+1}=\theta_{t}-\eta_{t}\tilde{g}_{t}$.
@@ 5.4. SNN Exploration Space
In the SNN exploration space as shown in Figure 1, we investigate two major learning algorithms: surrogate gradient-based learning  (Neftci et al., 2019) and evolutionary learning (Schuman et al., 2020b).
@@ 5.4.1. Surrogate Gradient based Learning:
Surrogate gradient-based learning algorithms are used to train Spiking Neural Networks (SNNs) by approximating the non-differentiable spiking function with a differentiable surrogate. This allows the application of gradient-based optimization techniques, such as backpropagation, to SNNs. By replacing the true gradient of the spike function with a smooth, differentiable approximation, these algorithms enable the effective training of SNNs. In our experiments, we employ the Arc-tangent  (Rajan et al., 2006) surrogate gradient algorithm through the snnTorch  (SNN-Torch Development, 2021) framework and the Spike-Timing-Dependent Plasticity (STDP)  (Caporale and Dan, 2008) learning rule using the LAVA (Intel Labs, 2023) framework by Intel.
@@
snnTorch is a deep learning framework specifically designed for SNNs built on PyTorch. For our experiments, we utilized the arc tangent learning algorithm within the framework. This algorithm approximates the non-differentiable spiking function with an arc tangent function. This framework was applied to MNIST, F-MNIST, Iris and Breast Cancer datasets to evaluate the resilience of SNNs against MIA.
In SNNs, encoding refers to the process of converting input data into spike trains that the network can process and it determines how input signals are represented as spikes, which regulates performance and behavior of the SNNs.
We explore the impact of different encoding schemes on privacy preserving properties of SNNs. The three encoding schemes  (Eshraghian, 2023) we investigated on snnTorch are:
@@
LAVA is an innovative open-source software framework designed for neuromorphic computing  (Roy et al., 2019). It provides a versatile and flexible environment for developing, simulating, and deploying neuromorphic applications. By abstracting hardware details through a process-based model, LAVA enables the creation of scalable and modular systems that can operate asynchronously across various neuromorphic and conventional hardware platforms. The framework supports rapid prototyping and detailed optimization, making it accessible to both researchers and developers interested in leveraging the unique capabilities of neuromorphic technologies. LAVA aligns particularly well with Intel’s Loihi chip  (Davies et al., 2018), a specialized neuromorphic processor. LAVA provides a seamless interface for developing applications that can efficiently run on Loihi. The LAVA framework also supports Spike-Timing-Dependent Plasticity (STDP), a biological learning rule that adjusts synaptic strengths based on the precise timing of spikes. This mechanism allows SNNs to learn temporal patterns and adapt to dynamic inputs. In our experiment, we used framework to evaluate the model’s resilience against privacy attack (MIA).
@@
MNIST Baseline 99.96% 99.21% 0.59 99.95% 99.22% 0.52
ResNet18 99.98% 99.57% 0.61 99.96% 99.51% 0.51
VGG16 99.67% 99.50% 0.61 99.58% 99.37% 0.51
F-MNIST Baseline 99.52% 92.77% 0.64 99.42% 92.44% 0.54
ResNet18 98.72% 93.66% 0.67 99.12% 91.79% 0.53
VGG16 97.14% 93.06% 0.62 96.53% 90.34% 0.55
CIFAR-10 Baseline 99.24% 73.20% 0.82 99.13% 72.99% 0.59
ResNet18 97.65% 90.81% 0.63 96.74% 86.85% 0.55
VGG16 97.45% 88.74% 0.67 79.77% 74.92% 0.53
CIFAR-100 Baseline 98.31% 42.46% 0.88 99.42% 39.92% 0.58
ResNet18 98.66% 70.92% 0.77 88.43% 60.65% 0.66
VGG16 81.61% 58.32% 0.70 60.53% 52.32% 0.58
Iris Baseline 100% 96.67% 0.77 97.04% 96.67% 0.57
Breast Cancer Baseline 99.54% 97.37% 0.65 99.62% 96.98% 0.55
Figure 5. Overfitting analysis in ANN and SNN models across datasets: F-MNIST, Breast Cancer, and CIFAR-10
@@ 5.4.2. Evolutionary Learning Algorithm:
Evolutionary algorithms, such as the Evolutionary Optimization for Neuromorphic Systems (EONS)  (Schuman et al., 2020b), are capable of enabling the rapid prototyping of SNN applications. These applications can be adapted to hardware constraints and various learning scenarios, including classification  (Schuman et al., 2020a, 2022a) and control tasks (Patton et al., 2021; Schuman et al., 2022b). EONS, implemented within the TENNLab framework, facilitates the co-design process between neuromorphic hardware and software, helping developers optimize neural network structures within the constraints of neuromorphic hardware. EONS interacts seamlessly with a wide variety of devices  (Schuman et al., 2020b), architectures  (Schuman et al., 2020c), and application  (Plank et al., 2018) without necessitating any modifications to its underlying algorithm. This approach, implemented within the TENNLab framework, utilizes an evolutionary algorithm to optimize neuromorphic network structure and weights. This process begins with generating a population of potential network solutions, which can be either randomly created or seeded with pre-existing networks  (Schuman et al., 2020b). The EONS process then evaluates these networks for fitness based on user-defined criteria and uses common genetic algorithm selection techniques such as tournament or fitness score to choose the most promising networks for reproduction. During reproduction, selected parent networks undergo crossover to swap segments of their structure and merging, which combines their entire structures into a single offspring. Mutation operations introduce random modifications to a network’s nodes and edges, facilitating diversity and adaptation in the population. These operations are performed on a generalized network representation consisting of nodes and edges, allowing flexibility and adaptability across different hardware implementations. The entire EONS cycle, designed to optimize network parameters and structure efficiently, repeats until achieving desired performance metrics.
@@
  • RISP  (Plank et al., 2022): A lightweight neuro processor employing an integrate-and-fire model with discrete time steps for accumulating and evaluating action potentials, suitable for networks with integer synaptic delays.
@@
Figure 6. ROC curves showing the impact of Rate, Latency, and Delta modulation encoding under MIA for MNIST, F-MNIST, CIFAR-10, Iris, and Breast Cancer datasets.

### v2

@@ 3. Methodology
This investigation examines the comparative privacy resilience of SNNs and ANNs through a systematic experimental framework comprising three distinct phases: (1) assessment of privacy vulnerabilities through MIA in both architectures, (2) analysis of SNN specific privacy characteristics across diverse algorithmic implementations, and (3) evaluation of privacy-utility trade offs through the implementation of DPSGD. The experimental methodology and interrelationships between architectural components are illustrated in Figure 1.
@@ 3.1. Comparison of Privacy Vulnerability between ANNs and SNNs
The comparative privacy risk assessment utilizes MIAs across equivalent ANN and SNN architectures, implementing convolutional baseline models, ResNet18, and VGG16 configurations. The experimental framework employs shadow models trained on labeled datasets (Figure 1, top-right) to emulate the target model characteristics, while attack models are developed to ascertain training set membership of individual data points. Privacy vulnerability is quantified through AUC metrics for both architectures, enabling systematic comparison of their susceptibility to MIAs (Figure 1, yellow block).
@@ 3.2. Algorithmic Exploration within the SNN Architecture
The second phase examines SNN privacy resilience across diverse algorithmic implementations through three distinct frameworks: Surrogate Gradient Algorithm, Evolutionary Algorithm, and Intel’s LAVA framework. The Surrogate Gradient implementation evaluates the baseline SNN model using three distinct encoding mechanisms from the snnTorch (Eshraghian, 2023) library: Delta, Latency, and Delta Modulation (Figure 1, left section, "SNN Exploration Space"). This analysis examines how encoding methods affect privacy vulnerability by correlating spike generation mechanisms with MIA efficacy.
The Evolutionary Algorithm implementation, utilizing the TennLab framework, modulates architectural parameters including population size, crossover rate, mutation rate, processor configurations, and encoder settings. This parameter space exploration evaluates how architectural variations influence SNN susceptibility to MIAs (Figure 1, green block).
@@ 3.3. Privacy-Utility Trade-off Analysis
The final experimental phase implements DPSGD across both ANN and SNN architectures ((Figure 1, bottom) to quantify the privacy-utility trade-offs. The DPSGD implementation incorporates Gaussian noise during gradient computation, establishing differential privacy guarantees while measuring the corresponding impact on model performance. The analysis spans convolutional, ResNet18, and VGG16 architectures, evaluating both model accuracy and resilience through attack AUC metrics under equivalent privacy constraints (Figure 1, blue block).
@@ 4.1. Dataset and Model Architecture
Network Variant Structure Parameters*
@@
The proposed method is evaluated on both image and tabular datasets. For image classification tasks, MNIST (LeCun et al., 2010) and Fashion-MNIST (Xiao et al., 2017), comprising 28×28 grayscale images across 10 classes, are utilized. CIFAR-10 (Recht et al., 2018) and CIFAR-100 (Sharma et al., 2018), containing 32×32 RGB images with 10 and 100 classes, respectively, are also included. Two tabular datasets, Iris (Omelina et al., 2021), consisting of 4 features and 3 classes, and Breast Cancer (Zwitter and Soklic, 1988), comprising 30 features and 2 classes, are used.
The experimental evaluation encompasses both image and tabular datasets. The image classification tasks utilize MNIST (LeCun et al., 2010) and Fashion-MNIST (Xiao et al., 2017), comprising 28×28 grayscale images across 10 classes, and CIFAR-10 (Recht et al., 2018) and CIFAR-100 (Sharma et al., 2018), containing 32×32 RGB images with 10 and 100 classes, respectively. The tabular datasets include Iris (Omelina et al., 2021) (4 features, 3 classes) and Breast Cancer (Zwitter and Soklic, 1988) (30 features, 2 classes).
The architectural implementations, as detailed in Table 1, are composed of three model configurations that are adapted for both ANN and SNN frameworks. The baseline architecture is implemented with dual convolutional layers for image processing and fully connected layers for tabular data analysis. ResNet18 is constructed with four basic blocks that incorporate group normalization and residual connections, while VGG16 is designed with 13 convolutional layers, progressively expanding from 64 to 512 channels, and culminating in three fully connected layers.
In the SNN variants, ReLU activations are replaced with leaky integrate-and-fire (LIF) neurons, and temporal processing mechanisms are incorporated. The implementation framework employs PyTorch (Ketkar et al., 2021) for ANN architectures, while snnTorch is used for the baseline model and SpikingJelly (Fang et al., 2023) is utilized for ResNet and VGG16 configurations.
@@ 4.2. MIA
[…]
The MIA framework employs a dual model architecture as illustrated in Figure 2. The experimental setup includes a ’target’ model and a ’shadow’ model, where we evaluate three different architectures: baseline model, ResNet18, and VGG16. The shadow model simulates the target model’s functionality by training on 80% of the corresponding dataset while maintaining architectural parity. For the attack phase, we employ a Support Vector Machine (SVM) as our attack model, chosen for its effectiveness as a binary classifier to determine whether a data point is a member or non-member of the training set. We evaluate the models by querying with their respective training and test sets, labeling the responses as ’IN’ if the data was in the model’s training set and ’OUT’ otherwise. To reduce classifier bias towards more frequent classes, we implement undersampling of the predominant class. The SVM attack model, trained on these labeled predictions from the shadow model, provides a quantified measure of the model’s susceptibility to MIA through its ability to accurately classify membership status.
@@ 4.3. DPSGD
[…]
The DPSGD implementation utilizes the Opacus Privacy Engine (Yousefpour et al., 2021), as illustrated in Figure 3. The DPSGD algorithm  1 begins by selecting a minibatch $L_{t}$ uniformly sampled from the set of indices $\{1,\dots,N\}$ with size $L$. For each data point $x_{i}$ in the minibatch, the gradient of the loss function $\mathcal{L}(\theta_{t},x_{i})$ with respect to the parameters $\theta_{t}$ is computed as $g_{t}(x_{i})$. These gradients are then clipped using an $L2$ norm operation, $\hat{g}_{t}(x_{i})=g_{t}(x_{i})/\max(1,\|g_{t}(x_{i})\|_{2}/C)$. Gaussian noise $\mathcal{N}(0,\sigma^{2}C^{2}I)$ is added to the average of the clipped gradients, with $\sigma$ being the noise scale, to ensure differential privacy. The noisy gradient $\tilde{g}_{t}$ is used to update the model parameters: $\theta_{t+1}=\theta_{t}-\eta_{t}\tilde{g}_{t}$.
@@ 4.4. SNN Exploration Space
In the SNN exploration space as shown in Figure 1, we investigate two major learning algorithms: surrogate gradient based learning  (Neftci et al., 2019) and evolutionary learning (Schuman et al., 2020b).
@@ 4.4.1. Surrogate Gradient based Learning:
Surrogate gradient methodologies facilitate SNN training through differentiable approximations of the non differentiable spiking activation function, enabling gradient based optimization techniques. These algorithms substitute the discrete spike function gradient with continuous, differentiable surrogates, permitting backpropagation based training while preserving the essential characteristics of spike-based computation. The experimental implementation employs two distinct approaches: the Arc tangent surrogate gradient algorithm  (Rajan et al., 2006) implemented via snnTorch  (SNN-Torch Development, 2021) framework, and the biologically-inspired Spike Timing Dependent Plasticity (STDP)  (Caporale and Dan, 2008) learning mechanism through Intel’s LAVA framework  (Intel Labs, 2023).
@@
The snnTorch framework, built upon PyTorch, provides specialized implementations for Spiking Neural Networks. Our experimental framework implements the arc tangent learning algorithm, which provides differentiable approximations of the spiking activation function. This implementation is evaluated across MNIST, F-MNIST, Iris, and Breast Cancer datasets to assess architectural resilience against membership inference attacks. In SNN architectures, the encoding mechanism determines the transformation of input signals into temporal spike patterns, fundamentally influencing the network’s computational characteristics and performance metrics. Our investigation examines the relationship between encoding methodologies and privacy preservation characteristics across three distinct encoding schemes  (Eshraghian, 2023):
@@
LAVA is an innovative open source software framework designed for neuromorphic computing  (Roy et al., 2019). It provides a versatile and flexible environment for developing, simulating, and deploying neuromorphic applications. By abstracting hardware details through a process-based model, LAVA enables the creation of scalable and modular systems that can operate asynchronously across various neuromorphic and conventional hardware platforms. The framework supports rapid prototyping and detailed optimization, making it accessible to both researchers and developers interested in leveraging the unique capabilities of neuromorphic technologies. LAVA aligns particularly well with Intel’s Loihi chip  (Davies et al., 2018), a specialized neuromorphic processor. LAVA provides a seamless interface for developing applications that can efficiently run on Loihi. The LAVA framework also supports Spike-Timing-Dependent Plasticity (STDP), a biological learning rule that adjusts synaptic strengths based on the precise timing of spikes. This mechanism allows SNNs to learn temporal patterns and adapt to dynamic inputs. In our experiment, we used framework to evaluate the model’s resilience against privacy attack (MIA).
@@
MNIST Baseline 99.96(±0.01)99.21(±0.03)0.59(±0.008) 99.95(±0.05)99.22(±0.02)0.52(±0.002)
ResNet18 99.98(±0.10)99.57(±0.04)0.61(±0.005) 99.96(±0.009)99.51(±0.06)0.51(±0.006)
VGG16 99.67(±0.05)99.50(±0.02)0.61(±0.007) 99.58(±0.12)99.37(±0.14)0.51(±0.004)
F-MNIST Baseline 99.52(±0.13)92.77(±0.20)0.64(±0.011 99.42(±0.32)92.44(±0.19)0.54(±0.008)
ResNet18 98.72(±0.21)93.66(±0.15)0.67(±0.004) 99.12(±0.09)91.79(±0.03)0.53(±0.003)
VGG16 97.14(±0.17)93.06(±0.20)0.62(±0.005) 96.53(±0.29)90.34(±0.19)0.55(±0.011)
CIFAR-10 Baseline 99.24(±0.08)73.20(±0.43)0.82(±0.095) 99.13(±0.45)72.99(±0.33)0.59(±0.005)
ResNet18 97.65(±0.51)90.81(±0.24)0.63(±0.021) 96.74(±0.44)86.85(±0.13)0.55(±0.011)
VGG16 97.45(±0.41)88.74(±0.29)0.67(±0.013) 79.77(±0.31)74.92(±0.20)0.53(±0.003)
CIFAR-100 Baseline 98.31(±0.19)42.46(±0.35)0.88(±0.016) 99.42(±0.59)39.92(±0.67)0.58(±0.019)
ResNet18 98.66(±0.23)70.92(±0.48)0.77(±0.009) 88.43(±0.27)60.65(±0.74)0.66(±0.002)
VGG16 81.61(±0.17)58.32(±0.52)0.70(±0.019) 60.53(±0.22)52.32(±1.36)0.58(±0.004)
Iris Baseline 100(±0.0)96.67(±2.34)0.77(±0.13) 97.04(±0.34)96.67(±0.0)0.57(±0.020)
Breast Cancer Baseline 99.54(±0.04)97.37(±0.24)0.65(±0.008) 99.62(±0.0)96.98(±0.17)0.55(±0.017)
Figure 5. Overfitting analysis in ANN and SNN models across datasets: (a) F-MNIST, (b) Breast Cancer, and (c) CIFAR-10
[…]
Figure 6. ROC curves showing the impact of Rate, Latency, and Delta modulation encoding under MIA for (a) MNIST, (b) F-MNIST, (c) CIFAR-10, (d) Iris, and (e) Breast Cancer datasets.
[…]
@@ 4.4.2. Evolutionary Learning Algorithm:
Evolutionary algorithms, such as the Evolutionary Optimization for Neuromorphic Systems (EONS)  (Schuman et al., 2020b), are capable of enabling the rapid prototyping of SNN applications. These applications can be adapted to hardware constraints and various learning scenarios, including classification  (Schuman et al., 2020a, 2022a) and control tasks (Patton et al., 2021; Schuman et al., 2022b). EONS, implemented within the TENNLab framework, facilitates the co-design process between neuromorphic hardware and software, helping developers optimize neural network structures within the constraints of neuromorphic hardware. EONS interacts seamlessly with a wide variety of devices  (Schuman et al., 2020b), architectures  (Schuman et al., 2020c), and application  (Plank et al., 2018) without necessitating any modifications to its underlying algorithm. This approach, implemented within the TENNLab framework, utilizes an evolutionary algorithm to optimize neuromorphic network structure and weights. This process begins with generating a population of potential network solutions, which can be either randomly created or seeded with pre-existing networks  (Schuman et al., 2020b). The EONS process then evaluates these networks for fitness based on user-defined criteria and uses common genetic algorithm selection techniques such as tournament or fitness score to choose the most promising networks for reproduction. During reproduction, selected parent networks undergo crossover to swap segments of their structure and merging, which combines their entire structures into a single offspring. Mutation operations introduce random modifications to a network’s nodes and edges, enhancing diversity and adaptation in the population. These operations are performed on a generalized network representation consisting of nodes and edges, allowing flexibility and adaptability across different hardware implementations. The entire EONS cycle, designed to optimize network parameters and structure efficiently, repeats until achieving desired performance metrics.
@@
  • RISP  (Plank et al., 2022): A lightweight neuro processor employing an integrate and fire model with discrete time steps for accumulating and evaluating action potentials, suitable for networks with integer synaptic delays.